# SigmaScheduling: Uncertainty-Informed Scheduling of Decision Points for Intelligent Mobile Health Interventions

### v1

@@ I Introduction
At predefined timepoints called “decision points,” just-in-time adaptive interventions (JITAIs) intelligently decide whether and how to provide mobile health support based on biobehavioral sensor or survey data [1]. A challenge for many JITAIs is that their effectiveness can depend on delivering interventions shortly before a target behavior, but identifying such pre-behavior time windows in real-time is often infeasible. For instance, while wrist-worn sensors can be used to detect eating [2], a reminder to take medication prior to a meal would be rendered ineffective if triggered only after the meal has already begun [3]. More commonly, the challenge arises due to practical constraints in implementation. For the “Oralytics” JITAI, which leverages a commercially available smart toothbrush and a companion app to promote oral self-care [4], behavioral data from the brush becomes available only once per night. As a result, decision points for a given day must be made in advance based on brushing patterns from previous days. Decision points are thus scheduled to anticipate – rather than respond to – forthcoming behavior.
@@
The status quo approach to scheduling decision points is to ask individuals to provide expected behavior times (e.g., brushing routines) prior to JITAI deployment and then schedule decision points the same fixed duration – for all users – before expected behavior times [5, 6, 7]. While this status quo may perform satisfactorily for individuals with relatively predictable routines, the status quo often falters for more variable behavior patterns, as shown in Fig. 1. Increasing the fixed duration from one to four hours could address the issue in Fig. 1B – but would unnecessarily increase the delay from decision points to brush times for individuals with more predictable routines (e.g., Fig. 1A). In settings where behavior predictability varies widely across the population, the one-size-fits-all status quo can squander valuable intervention opportunities. For example, nearly a quarter of participants in the first deployment of Oralytics had 70% or fewer of their brushing events preceded by a decision point.
This paper proposes and evaluates SigmaScheduling: a method to dynamically schedule decision points based on uncertainty in predicted behavior times. Our hypothesis is that SigmaScheduling will more effectively precede target behaviors with decision points by scheduling decision points further in advance when routines are less predictable and closer beforehand otherwise. To test this hypothesis, we apply SigmaScheduling to real-world data from 68 participants of a 10-week trial of Oralytics [4], using two approaches for brush time prediction and uncertainty quantification (UQ): one based on user-provided times and past errors and one based on online Bayesian machine learning (ML). Our contribution is to demonstrate that for both approaches, SigmaScheduling increases the proportion of participants whose brushing events are preceded by a decision point at least 70% of the time.
@@ II-A SigmaScheduling
The key design choice for SigmaScheduling is the parameter $c$. Selecting $c$ involves a tradeoff between “coverage,” the proportion of behavior times preceded by decision points (i.e., $\frac{1}{K}\cdot P$, where $P=\sum_{k=1}^{K}{\mathbbm{I}[t_{k}<t^{*}_{k}]}$), and the mean delay from decision points to behavior times (i.e., $\frac{1}{P}\sum_{k=1}^{K}{\mathbbm{I}[t_{k}<t^{*}_{k}]\cdot(t^{*}_{k}-t_{k})}$). In general, $c$ can vary across time (i.e., $\{c_{k}\}_{k=1}^{K}$) or individuals, but in this work, $c$ is fixed.
@@ II-D Oralytics Data Used for Evaluation
In this paper, we evaluate SigmaScheduling using brush time data from 68 participants in the first trial of Oralytics [4]. Of the 72 participants analyzed by Trella et al., we excluded one with less than a week of analyzable brush time data and three with corrupted data due to user-provided brush times between 0:00 and 04:00. Participants were adults at risk for dental disease who participated in a 70-day micro-randomized trial. Two decision points were scheduled per day, one for morning and one for evening oral self-care behavior. The trial was approved by the University of California, Los Angeles Institutional Review Board (IRB#21–001471), registered on ClinicalTrials.gov (NCT05624489), and lasted from Sept. 2023 to July 2024. See prior work for further details [4, 5].
Participants provided their typical brush times for morning and evening brushing on weekdays and weekends. Ground truth brush times, $\{t^{*}_{k}\}_{k=1}^{K}$, were sensed via smartbrush (Oral-B 8000). Only brush times, $t^{*}_{k}$, without prior intervention for the same $k$ were analyzed to focus on self-determined behavior times (decision points should precede these times). A mean of 53 $\pm$ 23 ($\pm$ SD) brush times were analyzed per participant.
@@ II-E Online Bayesian Linear Regression for Oralytics
Leave-one-participant-out cross validation was used to produce $N$ participant-specific Bayesian linear regression (BLR) models for brush time prediction and UQ. BLR was used to match the upcoming deployment of Oralytics. Gaussian priors were set for regression weights similar to prior work [4]. Prior means were set via a generalized estimating equation (GEE) fit to the $N-1$ participants’ training data: the GEE point estimate was used for statistically significant regressors; zero otherwise. Prior SDs were set by fitting a GEE model to each training participant’s data: the SD of point estimates was used for statistically significant regressors; half the SD otherwise. An inverse-gamma prior was used for the observation noise variance. The regressors were user-provided brush times; day of the week; morning or evening; the minimum, maximum, and coefficient of variation of past week brush times, most recent brush time; and number of no-brush days. For the held-out participant, priors were updated online with $\{t^{*}_{i}\}_{i=1}^{k-1}$ before predicting $\hat{t}_{k}$. The SD of the Gaussian posterior, $P(t^{*}_{k}|\hat{t}_{k})$, was used for $\sigma_{k}$ in (1), likening $c$ to a z-score.
@@ III Results
Fig. 3 compares status quo scheduling and SigmaScheduling via ROC-like curves described in II-F. Interpreting Fig. 3B as an example, if the goal is for participants to have at least 8/10 of their brush times preceded by decision points, status quo scheduling achieves this at a mean delay of $\sim$4.5 hours for $\sim$85of participants. For the same mean delay, SigmaScheduling achieves the desired coverage for $\sim$95of participants. Table I details the areas under the curves.
@@ IV Discussion
Using 68 participants’ real-world data from the first deployment of the Oralytics JITAI, we find that SigmaScheduling improves the likelihood that interventions are delivered at opportune moments, before the target behavior of tooth brushing, in over 7/10 of instances. The improvement stems from SigmaScheduling’s core principle of personalization: decision points are scheduled a longer duration before predicted brush times for participants with less predictable routines, while using shorter durations otherwise. In comparison, status quo scheduling uses the same duration for all participants, resulting in lead times that are often too long or too short.
SigmaScheduling’s advantage becomes more pronounced as desired coverage increases. However, under more lenient criteria such as when covering only 2/3 of brushing events per participant, status quo scheduling performs comparably or better. One of the primary reasons for this is that for participants with predictable routines, the majority of true brush times fall fairly close to predicted brush times. If it is acceptable to exclude the remaining minority (i.e., the threshold for desired coverage is lower), status quo scheduling suffices. Also, participants with irregular routines are the minority in the data collected. Status quo suffices when accepting a lower % of participants with desired coverage – in other words, when excluding participants with less predictable routines.
This work has limitations. Our evaluation of SigmaScheduling is limited to retrospective analysis within the oral self-care domain. Future studies are needed to deploy SigmaScheduling and for evaluation in domains such as dietary behavior [6, 7].
@@ V Conclusion
This paper introduces SigmaScheduling, a novel method for scheduling JITAI decision points that dynamically adapts the lead time between decision points and predicted behavior times to the uncertainty in predicted times. Applied to Oralytics data, SigmaScheduling more consistently schedules decision points prior to toothbrushing (i.e., during windows of receptivity [1]). Future work will evaluate SigmaScheduling in the upcoming deployment of Oralytics and apply SigmaScheduling to other domains such as dietary behavior.

### v2

@@ I Introduction
At predefined timepoints called “decision points,” just-in-time adaptive interventions (JITAIs) intelligently decide whether and how to provide mobile health support based on biobehavioral sensor or survey data [1]. A challenge for many JITAIs is that their effectiveness can depend on delivering interventions shortly before a target behavior, but identifying such pre-behavior time windows in real-time is often infeasible. For instance, while wrist-worn sensors can be used to detect eating [2], a reminder to take medication prior to a meal would be rendered ineffective if triggered only after the meal has already begun [3]. More commonly, the challenge arises due to practical constraints in implementation. For the “Oralytics” JITAI, which leverages a commercially available smart toothbrush and a companion app to promote oral self-care [4], data from the brush becomes available only once each night. As a result, decision points for a given day must be scheduled in advance based on prior brushing patterns to anticipate – rather than respond to – forthcoming behavior.
@@
The status quo approach to scheduling decision points is to ask individuals to provide expected behavior times prior to JITAI deployment and then schedule decision points the same fixed duration – for all users – before expected behavior times [5, 6, 7]. While this status quo may perform satisfactorily for individuals with relatively predictable routines, the status quo often falters for more variable behavior patterns, as shown in Fig. 1. Increasing the fixed duration from one to four hours could address the issue in Fig. 1B – but would unnecessarily increase the delay from decision points to brush times for more predictable routines (e.g., Fig. 1A). In settings where behavior predictability varies widely across the population, the one-size-fits-all status quo can squander valuable intervention opportunities. For example, nearly a quarter of participants in the first deployment of Oralytics had 70% or fewer of their brushing events preceded by a decision point.
This paper proposes and evaluates SigmaScheduling: a method to dynamically schedule decision points based on uncertainty in predicted behavior times. Our hypothesis is that SigmaScheduling will more effectively precede target behaviors with decision points by scheduling decision points earlier when routines are less predictable and closer beforehand otherwise. To test this hypothesis, we apply SigmaScheduling to real-world data from 68 participants of a 10-week trial of Oralytics [4], using (1) user-provided times and (2) online Bayesian machine learning (ML) for brush time prediction and uncertainty quantification (UQ). Our contribution is to demonstrate that for both approaches, SigmaScheduling increases the proportion of participants whose brushing events are preceded by a decision point at least 70% of the time.
@@ II-A SigmaScheduling
The key design choice for SigmaScheduling is the parameter $c$. Theoretically, $c$ represents a critical value and can be selected based on the desired probability that $t_{k}^{*}>t_{k}$. For example, if the posterior distribution $P(t_{k}^{*}|\hat{t}_{k})$ is modeled as a Gaussian distribution, then $c$ represents a z-score. In practice, selecting $c$ involves a tradeoff between “coverage,” the proportion of behavior times preceded by decision points (i.e., $\frac{1}{K}\cdot P$, where $P=\sum_{k=1}^{K}{\mathbbm{I}[t_{k}<t^{*}_{k}]}$), and the mean delay from decision points to behavior times (i.e., $\frac{1}{P}\sum_{k=1}^{K}{\mathbbm{I}[t_{k}<t^{*}_{k}]\cdot(t^{*}_{k}-t_{k})}$). In general, $c$ can vary across time (i.e., $\{c_{k}\}_{k=1}^{K}$) or individuals, but in this work, $c$ is fixed.
@@ II-D Oralytics Data Used for Evaluation
This paper evaluates SigmaScheduling using brush time data from 68 participants in the first trial of Oralytics [4]. Of the 72 participants analyzed by Trella et al., we excluded one with less than a week of analyzable brush times and three with corrupted data due to user-provided brush times between 0:00 and 04:00. Participants were adults at risk for dental disease. The 70-day micro-randomized trial had two daily decision points (morning and evening); was approved by the University of California, Los Angeles Institutional Review Board (IRB#21–001471) and registered on ClinicalTrials.gov (NCT05624489) [5]; and lasted Sept. 2023 to July 2024 [4].
Participants provided typical morning and evening brush times on weekdays and weekends. Ground truth brush times, $\{t^{*}_{k}\}_{k=1}^{K}$, were sensed via smartbrush (Oral-B 8000). Only brush times, $t^{*}_{k}$, without prior intervention for the same $k$ were analyzed to focus on self-determined behavior times, which decision points should be scheduled to precede. A mean of 53 $\pm$ 23 ($\pm$ SD) brush times were analyzed per participant.
@@ II-E Online Bayesian Linear Regression for Oralytics
Leave-one-participant-out cross validation was used to produce $N$ participant-specific Bayesian linear regression (BLR) models for brush time prediction and UQ. BLR was used to match the upcoming deployment of Oralytics. Gaussian priors for regression weights were set as in prior work [4]. Prior means were set via a generalized estimating equation (GEE) fit to the N1N-1 participants’ training data: the GEE point estimate was used for statistically significant regressors; zero otherwise. Prior SDs were set by fitting a GEE model to each training participant’s data: the SD of point estimates was used for statistically significant regressors; half the SD otherwise. An inverse-gamma prior was used for the observation noise variance. The regressors were user-provided brush times; day of the week; morning or evening; the minimum, maximum, and coefficient of variation of past week brush times, most recent brush time; and number of no-brush days. For the held-out participant, priors were updated online with $\{t^{*}_{i}\}_{i=1}^{k-1}$ before predicting $\hat{t}_{k}$ and producing Gaussian posterior, $P(t^{*}_{k}|\hat{t}_{k})$.
@@ III Results
Fig. 3 compares status quo scheduling and SigmaScheduling via ROC-like curves described in II-F. Interpreting Fig. 3B as an example, if the goal is for participants to have at least 8/10 of their brush times preceded by decision points, status quo scheduling achieves this at a mean delay of $\sim$4.5 hours for $\sim$85% of participants. For the same mean delay, SigmaScheduling achieves the desired coverage for $\sim$95% of participants. Table I details the areas under the curves.
@@ IV Discussion
Using 68 participants’ real-world data from the first deployment of the Oralytics JITAI, we find that SigmaScheduling improves the likelihood that interventions are delivered at opportune moments, before the target behavior of tooth brushing, in over 7/10 of instances. The improvement stems from SigmaScheduling’s core principle of personalization: decision points are scheduled a longer duration before predicted brush times for participants with less predictable routines, while using shorter durations otherwise. In comparison, status quo scheduling uses the same duration for all participants.
SigmaScheduling’s advantage becomes more pronounced as desired coverage increases. However, under more lenient criteria such as when covering only 2/3 of brushing events per participant, status quo scheduling performs comparably or better. One of the primary reasons for this is that for participants with predictable routines, the majority of true brush times fall fairly close to predicted brush times. In this dataset, participants with irregular routines were the minority. If it is acceptable to exclude the minority of participants with less predictable routines, (i.e., the threshold for desired coverage is lower), status quo scheduling suffices.
This work has limitations. Our evaluation of SigmaScheduling is limited to retrospective analysis within the oral self-care domain. Future studies are needed to deploy SigmaScheduling, extend its application to domains such as dietary behavior [6, 7], and evaluate the behavioral impact of improved coverage.
@@ V Conclusion
This paper introduces SigmaScheduling, a novel method for scheduling JITAI decision points that dynamically adapts the lead time between decision points and predicted behavior times to the uncertainty in predicted times. Applied to Oralytics data, SigmaScheduling more consistently schedules decision points prior to toothbrushing (i.e., during windows of receptivity [1]). Future work will evaluate SigmaScheduling in the upcoming clinical trial deployment of Oralytics and apply SigmaScheduling to other domains such as dietary behavior.